# Canonical formalism in the Schwarzschild spacetime

### v2

@@ Abstract
.3 Organization of this paper
This article is organized as follows:
  • In Sec. I, we review the general Hamiltonian formalism for the study of motion, up to linear order in the body’s spin, as presented in Paper I. It then applies this formalism to the Schwarzschild spacetime. In particular, Sec. I.1 includes a brief overview of the evolution equations for the particle’s multipoles along its worldline and the spin supplementary condition. Then, in Sec. I.2, we summarize a covariant Hamiltonian formulation for these equations, starting from the $14$D phase space $\mathcal{M}$ with degeneracies, to a $12$D one $\mathcal{N}$ non-degenerate but inconsistent physically, and then to a $10$D physically consistent one $\mathcal{P}$; see FIG. 1 for an illustration of this reduction process. Sec. I.3 contains all the explicit formulae required to apply the general setup to a Schwarzschild background, with particular emphasis on symmetries and related constants of motion.
  • In Sec. II, we explore the Hamiltonian formulation on the $12$D phase space $\mathcal{N}$. In order to achieve this objective, we propose a canonical transformation that generalises the conventional orbital elements for Schwarzschild geodesics dynamics to those applicable to the dynamics of a spinning particle. We call the resulting coordinates the (relativistic) Andoyer variables. They are presented in Sec. II.1 for the geodesic case, and in Sec. II.2 in the spinning case, respectively. The physical interpretation of the Andoyer variables is presented in Sec. II.3, while Sec. II.4 provides a comprehensive account of the resulting $12$D Hamiltonian system, which serves as the foundation for the investigation of $\mathcal{N}$. An “effective” Hamiltonian formulation of the $12$D dynamics, which is designed for a practical implementation, is also presented in Sec. II.5.
  • In Sec. III, the aim is to restrict the analysis from the $12$D phase space $\mathcal{N}$ to its $10$D sub-manifold $\mathcal{P}$ where the spin supplementary condition holds identically. This is done by first applying the theory of constrained Hamiltonian systems in Sec. III.1, allowing us to go from $\mathcal{N}$ to $\mathcal{P}$. Because the coordinates inherited from $\mathcal{N}$ are not canonical on $\mathcal{P}$ any more, Sec. III.2 is dedicated to finding such a canonical system.
  • In Sec. IV, we start in Sec. IV.1 with a summary of this paper’s main result: a canonical formulation of the spinning body’s dynamics on the $10$D physical phase space $\mathcal{P}$, and how it describes the system’s dynamics in Sec. IV.2. We then propose two applications of the formalism. First, because the system is integrable, we derive action-angle variables and the associated Hamiltonian frequencies in ready-to-use formulae in Sec. IV.3. Second, we derive a complete solution of the circular-orbit dynamics for generic spin configuration in Sec. IV.4, including an emphasis on innermost stable circular orbits.
Finally, some technical aspects of our treatment are relegated to appendices, including the convention for our notations in App. A, a thorough Hamiltonian investigation of Schwarzschild geodesics in App. B, and a summary of the Marck tetrad formalism traditionally used in the literature, in App. C. Appendix D contains some detailed calculations relative to the construction of canonical coordinates in Sec. III.
@@ I.1.1 Dipole approximation and Killing vectors
Now we can state what “choosing an SSC” means: it amounts to a choice of observer (i.e., a choice of $v^{a}$) for whom the mass dipole $D_{(v)}^{b}$ vanishes. For this observer, all the content of $S^{ab}$ can be obtained from $S_{(v)}^{a}$ because $D^{b}_{(v)}=0$ implies $S^{ab}=\varepsilon^{ab}_{\phantom{ab}cd}v_{c}S^{d}_{(v)}$ according to Eq. (10). However, this does not mean that the spin and mass dipole vectors defined in another frame are irrelevant for the analysis. In fact, it is essential in our framework to make a distinction between a frame defining the SSC and that to carry out the mathematical analysis.
@@ I.1.2 Linearization, Tulczyjew-Dixon SSC and Killing-Yano tensors
to emphasize its dependence on the TD SSC. Combining Eqs. (8) and (10) as well as the identity $\varepsilon^{abcd}\varepsilon_{afgh}=-6\,\delta^{[b}_{f}\delta^{\phantom{|}\!c}_{g}\delta^{d]}_{h}$ (cf. [75]), the TD spin vector is found to satisfy
@@
Mathematically, the algebraic-differential system made of Eqs. (7) and (11) is now well-posed: it has $13$ equations (i.e. $10$ ODEs and $3$ algebraic equations) for $13$ unknowns. Physically, however, it is not self-consistent. Indeed, the MPTD system defined by Eq. (7) is the truncation, at dipolar order, of a more general system that results from the application of a multipolar expansion in Refs. [27, 29, 30]. Yet, it is well-known that, at the quadrupolar level, quadratic-in-spin corrections arise to the right-hand side of these evolution equations. Therefore, working at the dipolar level means that we have implicitly omitted quadratic-in-spin terms in Eq. (7), by assumption. To be consistent, we must apply a truncation of Eq. (7) at linear-in-spin order; otherwise, the MPTD system becomes ambiguous due to the partial description of quadratic-in-spin effects. The TD SSC allows us to perform this necessary linearization. This procedure can be rigorously performed as an expansion in powers of the small parameter $d/\mathcal{R}\ll 1$ involved in the multipolar expansion of arbitrary extended objects in Refs. [27, 30], with $d$ the typical length scale of the body and $\mathcal{R}$ that of the background spacetime. Since the spin angular momentum of an extended object scales as $S_{\circ}\sim\mu d$, we have the following small parameter at hand
@@
In any case, neglecting non-linear terms in the $\epsilon$ expansion, Eqs. (7) and (11) lead us to the equations of interest in this paper, namely the linear-in-spin, MPTD $+$ TD SSC system:
@@
The Rüdiger invariants in Eqs. (16) are referred to as quasi-invariants in Ref. [80] in the sense that (i) they are only conserved at linear order in spin [i.e., $\nabla_{u}\mathfrak{K}$ and $\nabla_{u}\mathfrak{Q}$ are $O(\epsilon^{2}$)] and (ii) their conservation holds only under the TD SSC in Eq. (11). We, however, do not need to make this distinction between invariance and quasi-invariance because we adopt the point of view that only the linear system in Eq. (15) is of physical relevance at dipolar order.
@@ I.2 Formulation as a covariant Hamiltonian system
Figure 1: The relevant phase space and its sub-manifolds for the covariant Hamiltonian formulation of the linear-in-spin system. We use different sub-manifolds to lift the degeneracies associated with the existence of Casimir invariants $(\mathcal{M}\rightarrow\mathcal{N})$ and correctly implement the TD SSC $(\mathcal{N}\rightarrow\mathcal{T}\rightarrow\mathcal{P})$. Geodesics dynamics are recovered in the non-spinning limit, spanning the 8D phase space $\mathcal{G}$.
@@ I.2.2 The $14$D Poisson manifold $\mathcal{M}$
The variables $(x^{\alpha},p_{\alpha},S^{\alpha\beta})$ or the variables $y_{\mathcal{M}}:=(x^{\alpha},\pi_{\alpha},S^{I},D^{I})$ are two different charts for the same $14$D phase space, denoted $\mathcal{M}$ [cf. TABLE 2 in Appendix. A]. Using the latter variables $y_{\mathcal{M}}$, which are more useful for practical calculations, the Poisson brackets for two arbitrary functions $F(y_{\mathcal{M}})$ and $G(y_{\mathcal{M}})$ on $\mathcal{M}$ are defined by [cf. Sec. II of Paper I]
@@
The Poisson brackets defined in Eqs. (21) are degenerate, i.e., the antisymmetric matrix $\Lambda^{ij}(y_{\mathcal{M}})$ has (non-maximal) rank $12$. The difference between $(\mathcal{M})$ and rank$(\Lambda)$ is precisely the number of Casimir invariants that exists for the brackets. These $14-12=2$ Casimir invariants, which we denote by $\mathcal{C}_{\circ}$ and $\mathcal{C}_{\star}$, have the unique (and defining) property that their Poisson bracket with any other function on $\mathcal{M}$ identically vanishes. In our case, these Casimirs are given by (cf Sec. II.D of Paper I)
@@ I.2.3 The $12$D symplectic leaves $\mathcal{N}$
The 14D phase space $\mathcal{M}$ is naturally foliated by 12D sub-manifolds, denoted $\mathcal{N}$, cf. Fig. 1. They are defined as the level sets of the two Casimir invariants defined in Eq. (22), and called symplectic leaves. Comparing the (phase-space) definitions in Eq. (22) with the (spacetime) spin norms $S_{\circ}$ and $S_{\star}$ in Eq. (13) readily reveals that, numerically, $\mathcal{C}_{\circ}=S_{\circ}^{2}$ and $\mathcal{C}_{\star}=S_{\star}^{2}$, hence the notation.
@@
Before leaving this subsection, it is very important to appreciate the distinction between the role played by the body’s mass $\mu$ and that by the spin norm $S_{\circ}$ in our Hamiltonian formalism, even though from the point of view of the MPTD equations, they have the same status. The constancy of $\mu$ along solutions is a consequence of our Hamiltonian in Eq. (27) which is autonomous (independent of $\bar{\tau}$). Inserting Eqs. (18a) into Eq. (27), and making the substitution in the definition of the dynamical mass $\mu$ in Eq. (8), we find that this “on-shell” value of $H_{\mathcal{N}}$ along trajectories is numerically equal to $-\mu^{2}/2$. In other words, the mass $\mu$ is conserved because it is a scalar field on $\mathcal{N}$ corresponding to a first integral of the system (we refer to Sec. IV of Paper I for a reminder about first integrals and other classical notions of Hamiltonian mechanics). On the other hand, the spin norm $S_{\circ}$ is a constant because as a result of the symplectic structure of $\mathcal{N}$ itself; it is the property of the phase space geometry. It would be constant for any other Hamiltonian, and may as well be considered as an external parameter on $\mathcal{N}$. In particular, it is not a first integral, as it is not a scalar field on $\mathcal{N}$ in the first place.
@@ I.2.4 The $10$D-Physical phase space $\mathcal{P}$
The issue here is that the TD SSC in Eq. (11) cannot be recovered from the linear-in-spin system in Eq. (15), even though the latter was originally derived with the help of the TD SSC in Sec. I.1. It is a one way procedure, and indeed Eqs. (15) only imply $\nabla_{u}(p_{a}S^{ab})=0$, not the TD SSC itself. This is not a problem when working directly at the level of the algebraic-differential system defined by Eqs. (15)-(11), i.e., outside Hamiltonian mechanics. Setting the TD SSC at some point on $\mathscr{L}$ as an initial condition, the parallel transport $\nabla_{u}(p_{a}S^{ab})=0$ always implies $p_{a}S^{ab}=0$ throughout $\mathscr{L}$ by linearity.
@@
For the purpose of later computations, it is useful to introduce the required constraints in terms of the tetrad components of the TD SSC $C^{A}:=\pi_{A}S^{AB}$, where $\pi^{I}=\eta^{IJ}(e_{J})^{\alpha}\pi_{\alpha}$ is the tetrad component of $\pi_{\alpha}$ and we use the relation $\pi_{a}=p_{a}+O(\epsilon)$ in Eq. (18a). Making use of Eq. (19), we then adopt the first two components of $C^{A}$ as the constraints in the form
@@
where $\pi_{I}$ are functions of $(x^{\alpha},\pi_{\alpha})$ and $(S^{I},D^{I})$ are given by Eq. (28). We note that any other pair $(C^{A},C^{B})$ of constraints from the TD SSC (or any combinations thereof), for fixed $A$ and $B$ in $0,\ldots,3$, can also be imposed as constraints. The reason for our specific choice in Eq. (30) will become clear when we concretely apply it to the Schwarzschild case in subsequent developments. The restriction from $\mathcal{N}$ to $\mathcal{P}$ via the TD SSC sub-manifold $\mathcal{T}\subset\mathcal{M}$ is best summarized in Fig. 1.

### v3

@@ Abstract
.3 Organisation of this paper and guide to the reader
We conclude this introduction with an overview of the subsequent sections of this paper a nd a guide for the reader, who may be interested in a specific subset of the results presented in our work.
In Sec. I, we shall give various preliminaries concerning the covariant Hamiltonian framework of the spinning secondary with a background Schwarzschild metric, up to linear order in the secondary spin. This section is aligned with the contents presented in Paper I [1]. Subsequently, in Sec. II we shall introduce the canonical variables on the $12$D phase space $\mathcal{N}$, namely the (relativistic) Andoyer variables, which make use of the spherical symmetry of Schwarzschild spacetime. The main technical details of the reduction from the $12$D phase space $\mathcal{N}$ to its $10$D physical sub-manifold $\mathcal{P}$, where the spin supplementary condition is satisfied identically, are provided in Sec. III. We finally present in Sec. IV the Hamiltonian system on the $10$D physical phase space $\mathcal{P}$ with the canonical coordinates, where the action-angle variables and Hamiltonian frequencies are also obtained. Some technical aspects of our treatment are relegated to appendices, including a thorough Hamiltonian investigation of Schwarzschild geodesics in App. A, and a summary of the Marck tetrad formalism in Ref. [55] which has traditionally been used in the literature, in App. B. Appendix C contains some detailed calculations relative to the construction of canonical coordinates in Sec. III.
The covariant Hamiltonian framework, as outlined in Sec I, is a foundational element of our series of works. In particular, Sec. I.3 contains all the explicit formulae required to apply this general framework to a specific Schwarzschild background, with an emphasis on the role of symmetries and related constants of motion. Therefore, we advise all our readers to familiarise themselves with this section, which sets out the conventions, notations and invariants of motion defined on Schwarzschild. However, the reader who have already studied our previous works may choose to skip this section.
The reader only interested in results concerning the newly defined canonical coordinate, i.e., the relativistic Andoyer variables on the $12$D symplectic manifold $\mathcal{N}$, need only read on Sec. II, particularly Sec. II.4 for a comprehensive account of the resulting $12$D Andoyer Hamiltonian system and Sec. II.5 for its practical ‘Hill’ formulation.
Conversely, those with only a passing interest in results pertaining to the canonical formulation on the $10$D physical manifold $\mathcal{P}$, need only refer to Sec. IV for details. Those seeking solely the practical results, such as the action-angle variables and Hamilton frequencies, are advised to turn directly to Sec. IV.3.
.4 Conventions and notations
[…]
Table 2: List of phase space coordinates used in this paper. The term “canonical” is used to indicate that the coordinate verifies the canonical Poisson matrix. The equation numbers indicate the locations within this paper where each coordinate is defined.
Manifold Coordinate set     Canonical Definition
14D-Poisson $(x^{\alpha},p_{\beta},S^{\gamma\delta})$ No Paper I
manifold $\mathcal{M}$ $(x^{\alpha},\pi_{\beta},S^{I},D^{I})$ No Paper I [Eq. (20)]
12D-Symplectic $y_{\mathcal{N}}:=(t,r,\theta,\phi,\sigma,\zeta,\pi_{t},\pi_{r},\pi_{\theta},\pi_{\phi},\pi_{\sigma},\pi_{\zeta})$ Yes Paper I [Eq. (29)]
manifold $\mathcal{N}$ $y_{\mathcal{N}}^{\textrm{A}}:=(t,r,\nu,\ell,s,\zeta,\pi_{t},\pi_{r},\pi_{\nu},\pi_{\ell},\pi_{s},\pi_{\zeta})$ Yes Andoyer [Eq. (66)]
10D-Physical $y_{\mathcal{P}}^{\textrm{C}}:=(t,r,\nu,\ell,\Gamma,\pi_{t},\pi_{r},\pi_{\nu},\pi_{\ell},\pi_{\varpi})$ No Complete [Eq. (80)]
phase space $\mathcal{P}$ $y_{\mathcal{P}}^{\textrm{F}}:=(\tilde{t},r,\nu,\tilde{\ell},{\varpi},\pi_{t},\tilde{\pi}_{r},\pi_{\nu},\pi_{\ell},\pi_{\varpi})$ Yes Final [Eq. (115)]
The symbol $O(\epsilon^{N})$ indicates an expression of order $N$ in the small parameter $\epsilon=S_{\circ}/\mu M$ [cf. Eq. (15)]. Unless otherwise specified, we work consistently at order $O(\epsilon)$, with terms of order $O(\epsilon^{n})$ with $n\geq 2$ being neglected. Furthermore, such terms are most often omitted from equations in order to enhance readability.
The various coordinate systems employed in the different phase spaces [cf. FIG. 1] described in the main body of the paper are summarized in Table 2.
@@ I.1.1 Dipole approximation and Killing vectors
Now we can state what “choosing an SSC” means: it amounts to a choice of observer (i.e., a choice of $v^{a}$) for whom the mass dipole $D_{(v)}^{b}$ vanishes. For this observer, all the content of $S^{ab}$ can be obtained from $S_{(v)}^{a}$ because $D^{b}_{(v)}=0$ implies $S^{ab}=\varepsilon^{ab}_{\phantom{ab}cd}v^{c}S^{d}_{(v)}$ according to Eq. (11). However, this does not mean that the spin and mass dipole vectors defined in another frame are irrelevant for the analysis. In fact, it is essential in our framework to make a distinction between a frame defining the SSC and that to carry out the mathematical analysis.
@@ I.1.2 Linearization, Tulczyjew-Dixon SSC and Killing-Yano tensors
to emphasize its dependence on the TD SSC. Combining Eqs. (9) and (11) as well as the identity $\varepsilon^{abcd}\varepsilon_{afgh}=-6\,\delta^{[b}_{f}\delta^{\phantom{|}\!c}_{g}\delta^{d]}_{h}$ [cf. a text book [97]], the TD spin vector is found to satisfy
@@
Mathematically, the algebraic-differential system made of Eqs. (8) and (12) is now well-posed: it has $13$ equations (i.e. $10$ ODEs and $3$ algebraic equations) for $13$ unknowns. Physically, however, it is not self-consistent. Indeed, the MPTD system defined by Eq. (8) is the truncation, at dipolar order, of a more general system that results from the application of a multipolar expansion in Refs. [47, 49, 50]. Yet, it is well-known that, at the quadrupolar level, quadratic-in-spin corrections arise to the right-hand side of these evolution equations. Therefore, working at the dipolar level means that we have implicitly omitted quadratic-in-spin terms in Eq. (8), by assumption. To be consistent, we must apply a truncation of Eq. (8) at linear-in-spin order; otherwise, the MPTD system becomes ambiguous due to the partial description of quadratic-in-spin effects. The TD SSC allows us to perform this necessary linearization. This procedure can be rigorously performed as an expansion in powers of the small parameter $d/\mathcal{R}\ll 1$ involved in the multipolar expansion of arbitrary extended objects in Refs. [47, 50], with $d$ the typical length scale of the body and $\mathcal{R}$ that of the background spacetime. Since the spin angular momentum of an extended object scales as $S_{\circ}\sim\mu d$, we have the following small parameter at hand [106]
@@
In any case, neglecting non-linear terms in the $\epsilon$ expansion, Eqs. (8) and (12) lead us to the equations of interest in this paper, namely the linear-in-spin, MPTD system:
@@
The Rüdiger invariants in Eqs. (17) are referred to as quasi-invariants in Ref. [111] in the sense that (i) they are only conserved at linear order in spin [i.e., $\nabla_{u}\mathfrak{K}$ and $\nabla_{u}\mathfrak{Q}$ are $O(\epsilon^{2}$)] and (ii) their conservation holds only under the TD SSC in Eq. (12).
@@ I.2.2 The $14$D Poisson manifold $\mathcal{M}$
The variables $(x^{\alpha},p_{\alpha},S^{\alpha\beta})$ or the variables $y_{\mathcal{M}}:=(x^{\alpha},\pi_{\alpha},S^{I},D^{I})$ are two different charts for the same $14$D phase space, denoted $\mathcal{M}$ [cf. TABLE 2 in Sec. .4]. Using the latter variables $y_{\mathcal{M}}$, which are more useful for practical calculations, the Poisson brackets for two arbitrary functions $F(y_{\mathcal{M}})$ and $G(y_{\mathcal{M}})$ on $\mathcal{M}$ are defined by [cf. Sec. II of Paper I]
@@
The Poisson brackets defined in Eqs. (22) are degenerate, i.e., the antisymmetric matrix $\Lambda^{ij}(y_{\mathcal{M}})$ has (non-maximal) rank $12$. The difference between $(\mathcal{M})$ and rank$(\Lambda)$ is precisely the number of Casimir invariants that exists for the brackets. These $14-12=2$ Casimir invariants, which we denote by $\mathcal{C}_{\circ}$ and $\mathcal{C}_{\star}$, have the unique (and defining) property that their Poisson bracket with any other function on $\mathcal{M}$ identically vanishes. In our case, these Casimirs are given by [cf. Sec. II.D of Paper I]
@@ I.2.3 The $12$D symplectic leaves $\mathcal{N}$
The 14D phase space $\mathcal{M}$ is naturally foliated by 12D sub-manifolds, denoted $\mathcal{N}$ [cf. Fig. 1]. They are defined as the level sets of the two Casimir invariants defined in Eq. (23), and called symplectic leaves. Comparing the (phase-space) definitions in Eq. (23) with the (spacetime) spin norms $S_{\circ}$ and $S_{\star}$ in Eq. (14) readily reveals that, numerically, $\mathcal{C}_{\circ}=S_{\circ}^{2}$ and $\mathcal{C}_{\star}=S_{\star}^{2}$, hence the notation.
@@
Before leaving this subsection, it is very important to appreciate the distinction between the role played by the body’s mass $\mu$ and that by the spin norm $S_{\circ}$ in our Hamiltonian formalism, even though from the point of view of the MPTD equations, they have the same status. The constancy of $\mu$ along solutions is a consequence of our Hamiltonian in Eq. (28) which is autonomous (independent of $\bar{\tau}$). Inserting Eqs. (19a) into Eq. (28), and making the substitution in the definition of the dynamical mass $\mu$ in Eq. (9), we find that this “on-shell” value of $H_{\mathcal{N}}$ along trajectories is numerically equal to $-\mu^{2}/2$. In other words, the mass $\mu$ is conserved because it is a scalar field on $\mathcal{N}$ corresponding to a first integral of the system (we refer to Sec. IV of Paper I for a reminder about first integrals and other classical notions of Hamiltonian mechanics). On the other hand, the spin norm $S_{\circ}$ is a constant because as a result of the symplectic structure of $\mathcal{N}$ itself; it is the property of the phase space geometry. It would be constant for any other Hamiltonian, and may as well be considered as an external parameter on $\mathcal{N}$. In particular, it is not a first integral, as it is not a scalar field on $\mathcal{N}$ in the first place, and therefore should be counted in the list of integrals of motion (to verify integrability, for example). We note that $S_{\circ}$ being a Casimir of the Poisson bracket also happens for the Newton-Wigner SSC, for which the algebra is SO$(3)$ [cf. [89]].
@@ I.2.4 The $10$D-Physical phase space $\mathcal{P}$
The issue here is that the TD SSC in Eq. (12) cannot be recovered from the linear-in-spin system in Eq. (16), even though the latter was originally derived with the help of the TD SSC in Sec. I.1. It is a one way procedure, and indeed Eqs. (16) only imply $\nabla_{u}(p_{a}S^{ab})=0$, not the TD SSC itself. This is not a problem when working directly at the level of the algebraic-differential system defined by Eqs. (16)-(12), i.e., outside Hamiltonian mechanics. Suppose that the TD SSC $p_{a}S^{ab}=0$ is imposed at a point on the worldline $\mathscr{L}$. Then, the parallel transport $\nabla_{u}(p_{a}S^{ab})=0$ of the SSC (as implied by Eqs. (16) up to quadratic corrections), guarantees that the TD SSC is maintained throughout the worldline.
@@
For the purpose of later computations, it is useful to introduce the required constraints in terms of the tetrad components of the TD SSC $C^{A}:=\pi_{A}S^{AB}$, where $\pi^{I}=\eta^{IJ}(e_{J})^{\alpha}\pi_{\alpha}$ is the tetrad component of $\pi_{\alpha}$. Making use of Eq. (20), we then adopt the first two components of $C^{A}$ as the constraints in the form
@@
where $\pi_{I}$ are functions of $(x^{\alpha},\pi_{\alpha})$ and $(S^{I},D^{I})$ are given by Eq. (29). We also used the approximation for the momentum: $p_{a}=\pi_{a}+O(\epsilon)$. We note that any other pair $(C^{A},C^{B})$ of constraints from the TD SSC (or any combinations thereof), for fixed $A$ and $B$ in $0,\ldots,3$, can also be imposed as constraints. The reason for our specific choice in Eq. (31) will become clear when we concretely apply it to the Schwarzschild case in subsequent developments. The restriction from $\mathcal{N}$ to $\mathcal{P}$ via the TD SSC sub-manifold $\mathcal{T}\subset\mathcal{M}$ is best summarized in Fig. 1.